# Equatorial periodic orbits and gravitational waveforms in a black hole free of Cauchy horizon

### v1

@@ I Introduction
For EMRIs, the mass ratio between the two objects is highly unequal, with the smaller object rapidly orbiting the supermassive black hole. The orbit gradually decays over time, and GWs are emitted during this process. Due to the extremely small fraction of energy carried away by the lower-mass object’s orbital motion compared to the system’s total energy, the inspiral timescale of the smaller-mass object around the supermassive black hole can extend over several years. During this period, the object can complete approximately $10^{3}\sim 10^{5}$ orbital cycles. The prolonged duration of the EMRI signals not only facilitates the accumulation of a higher signal-to-noise ratio Baiblack holeav:2021 , but also provides an extended observation window for precise parameter estimation in GW physics, thereby highlighting the significant scientific value of GWs emitted from EMRIs. From a fundamental physics perspective, since the GWs of EMRIs vary depending on the underlying gravitational theory, detecting these waveforms provides an approach to validate or challenge current models of gravity Amaro-Seoane:2019umn ; Liang:2022gdk . Thus, EMRIs offer an opportunity to test general relativity and probe modified theories of gravity. From an astronomical perspective, the detection of EMRIs can be used to accurately measure the intrinsic parameters of supermassive black holes, such as their mass, spin, and quadrupole moment Amaro-Seoane:2007osp . Furthermore, the presence of dark matter surrounding a supermassive black hole could influence the motion of the compact object through dynamical friction, potentially altering the EMRI waveform. Therefore, by analyzing EMRI signals, we can also investigate the existence of dark matter around supermassive black holes Yue:2017iwc ; Yue:2018vtk ; Duque:2023seg . In the field of cosmology, EMRIs serve as standard sirens, providing precise constraints on cosmological parameters Schutz:1986gp . Predictions for the occurrence rate of EMRI vary for different theoretical models, and most estimates suggest that future space-based GW detectors, such as LISA Danzmann:1997hm ; Schutz:1999xj ; Gair:2004iv ; LISA:2017pwj ; Maselli:2021men , Taiji Hu:2017mde , and Tianqin TianQin:2015yph , will be capable of detecting several, or even a large number of EMRI events Babak:2017tow ; Amaro-Seoane:2012lgq .
@@ II.1 The construction of Black hole without Cauchy horizon
It should be noted that the metric generated by the mass function (4) is not globally defined but is confined to the region $r\leq r_{s}$. The total mass $M$ of the system within $r<r_{s}$ is defined as $M\equiv m(r_{s})=r_{s}/2$, independent of the parameters $k$ and $n$. Therefore, the event horizon is equivalent to the Schwarzschild radius, with $r_{s}=2M$. In addition, different values of the parameters $k$ and $n$ can give rise to a wide range of interesting and diverse physical environments. For instance, when $k=1$, the metric derived from the mass function (4) satisfies the strong energy condition, while for $k>6$, it satisfies the dominant energy condition Ovalle:2023vvu . In Ref. Ovalle:2023vvu , considering the presence of a de Sitter core in regular black holes, only the case with $n=2$ is examined. Specifically, for ${k=1,n=2}$, it ensures the integrability of the scalar curvature $R=-4\Lambda+4/r^{2}$, meaning that the curvature remains finite. In the case of $n=2$, the parameter $k$ determines the degree of deformation of spacetime inside the object $(r<r_{s})$. When $k>1$, the deviation of the internal spacetime from anti-de Sitter space can be measured. Specifically, when $k$ takes integer values, the internal spacetime can be interpreted as a finite superposition of multiple anti-de Sitter spaces. This also serves as an important manifestation of the linearity of the Kerr-Schild mass function, which is given in Eq. (16) of Ref. Ovalle:2023vvu .
@@ II.2 The properties of black hole without Cauchy horizon
In Fig. 1(b), the behaviors of the connection point $r_{s}$ and the black hole horizon with varying hair parameter $\ell$ are shown. The black dot-dashed lines represent the two branches of solutions to Eq. (3), where the inner and outer mass functions correspond to $m(r)$ in Eq. (4) and $\tilde{m}(r)$ in Eq. (6), respectively. The gray dot-dashed line denotes the inner solution $r_{s_{(i)}}$, while the dark black dot-dashed line represents the outer solution $r_{s_{(o)}}$, which is also the horizon derived from $f(r)=0$ using the inner mass function $m(r)$. The purple lines correspond to the two branches of Eq. (8): the external dark purple solid line represents the event horizon $r_{h_{(o)}}$ (same as Eq. (9)), and the internal light purple dashed line represents the discarded “Cauchy horizon” $r_{h_{(i)}}$. The cyan dotted line marks the event horizon constructed by the inner and outer mass functions, coinciding with $r_{s_{(o)}}$. Outside this region lies the true exterior of the black hole. Here, the total mass $M$ is contained within $r_{s}$, but for $21\leq\ell/\mathcal{M}<$ 2, part of $\mathcal{M}$ is distributed on the surface $r_{s}$.
@@ III PERIODIC ORBITS
To illustrate the behavior of $q$, we present $q$ versus particle energy $E$ and angular momentum $L/\mathcal{M}$ in Fig. 6. In Figs. 6(a) and 6(b), the variations of the rational number $q$ with particle energy are shown. It is found that $q$ increases slowly with energy $E$ initially and then increases rapidly after reaching the maximum energy. In Fig. 6(a), the variation of $q$ with $\epsilon$ is shown for the case where the hair parameter $\ell/\mathcal{M}=1$ and $\epsilon$ takes different values. It is observed that $\epsilon$ has a significant impact on the distribution of $q-E$. In Fig. 6(b), we fix $\epsilon=0.5$ and explore the influence of different $\ell/\mathcal{M}$ on $q$. The results show that the closer the system is to the Schwarzschild case, the larger the maximum energy corresponding to $q$. On the other hand, in Figs. 6(c) and 6(d), the behaviors of $q$ as a function of the $L/\mathcal{M}$ are shown. It is found that $q$ decreases as the orbital angular momentum increases.

### v2

@@ I Introduction
For EMRIs, the mass ratio between the two objects is highly unequal, with the smaller object rapidly orbiting the supermassive black hole. The orbit gradually decays over time, and GWs are emitted during this process. Due to the extremely small fraction of energy carried away by the lower-mass object’s orbital motion compared to the system’s total energy, the inspiral timescale of the smaller-mass object around the supermassive black hole can extend over several years. During this period, the object can complete approximately $10^{3}\sim 10^{5}$ orbital cycles. The prolonged duration of the EMRI signals not only facilitates the accumulation of a higher signal-to-noise ratio Baiblack holeav:2021 , but also provides an extended observation window for precise parameter estimation in GW physics, thereby highlighting the significant scientific value of GWs emitted from EMRIs. From a fundamental physics perspective, since the GWs of EMRIs vary depending on the underlying gravitational theory, detecting these waveforms provides an approach to validate or challenge current models of gravity Amaro-Seoane:2019umn ; Liang:2022gdk ; AbhishekChowdhuri:2023gvu ; Kumar:2024utz . Thus, EMRIs offer an opportunity to test general relativity and probe modified theories of gravity. From an astronomical perspective, the detection of EMRIs can be used to accurately measure the intrinsic parameters of supermassive black holes, such as their mass, spin, and quadrupole moment Amaro-Seoane:2007osp . Furthermore, the presence of dark matter surrounding a supermassive black hole could influence the motion of the compact object through dynamical friction, potentially altering the EMRI waveform. Therefore, by analyzing EMRI signals, we can also investigate the existence of dark matter around supermassive black holes Yue:2017iwc ; Yue:2018vtk ; Duque:2023seg ; Rahman:2023sof . In the field of cosmology, EMRIs serve as standard sirens, providing precise constraints on cosmological parameters Schutz:1986gp . Predictions for the occurrence rate of EMRI vary for different theoretical models, and most estimates suggest that future space-based GW detectors, such as LISA Danzmann:1997hm ; Schutz:1999xj ; Gair:2004iv ; LISA:2017pwj ; Maselli:2021men ; Ghosh2024 , Taiji Hu:2017mde , and Tianqin TianQin:2015yph , will be capable of detecting several, or even a large number of EMRI events Babak:2017tow ; Amaro-Seoane:2012lgq .
@@ II.1 The construction of Black hole without Cauchy horizon
It should be noted that the metric generated by the mass function (4) is not globally defined but is confined to the region $r\leq r_{s}$. The total mass $M$ of the system within $r<r_{s}$ is defined as $M\equiv m(r_{s})=r_{s}/2$, independent of the parameters $k$ and $n$. Therefore, the event horizon is equivalent to the Schwarzschild radius, with $r_{s}=2M$. In addition, different values of the parameters $k$ and $n$ can give rise to a wide range of interesting and diverse physical environments. For instance, when $k=1$, the metric derived from the mass function (4) satisfies the strong energy condition, while for $k>6$, it satisfies the dominant energy condition Ovalle:2023vvu . In Ref. Ovalle:2023vvu , considering the presence of a de Sitter core in regular black holes, only the case with $n=2$ is examined. Specifically, for k=1,n=2{k=1,n=2}, it ensures the integrability of the scalar curvature $R=-4\Lambda+4/r^{2}$, meaning that the curvature remains finite. In the case of $n=2$, the parameter $k$ determines the degree of deformation of spacetime inside the object $(r<r_{s})$. When $k>1$, the deviation of the internal spacetime from anti-de Sitter space can be measured. Specifically, when $k$ takes integer values, the internal spacetime can be interpreted as a finite superposition of multiple anti-de Sitter spaces. This also serves as an important manifestation of the linearity of the Kerr-Schild mass function, which is given in Eq. (16) of Ref. Ovalle:2023vvu .
@@ II.2 The properties of black hole without Cauchy horizon
In Fig. 1(b), the behaviors of the connection point $r_{s}$ and the black hole horizon with varying hair parameter $\ell$ are shown. The black dot-dashed lines represent the two branches of solutions to Eq. (3), where the inner and outer mass functions correspond to $m(r)$ in Eq. (4) and $\tilde{m}(r)$ in Eq. (6), respectively. The gray dot-dashed line denotes the inner solution $r_{s_{(i)}}$, while the dark black dot-dashed line represents the outer solution $r_{s_{(o)}}$, which is also the horizon derived from $f(r)=0$ using the inner mass function $m(r)$. The purple lines correspond to the two branches of Eq. (8): the external dark purple solid line represents the event horizon $r_{h_{(o)}}$ (same as Eq. (9)), and the internal light purple dashed line represents the discarded “Cauchy horizon” $r_{h_{(i)}}$. The cyan dotted line marks the event horizon constructed by the inner and outer mass functions, coinciding with $r_{s_{(o)}}$. Outside this region lies the true exterior of the black hole. Here, the total mass $M$ is contained within $r_{s}$, but for $1\leq\ell/\mathcal{M}<2$, part of $\mathcal{M}$ is distributed on the surface $r_{s}$.
@@ III PERIODIC ORBITS
To illustrate the behavior of $q$, we present $q$ versus particle energy $E$ and angular momentum $L/\mathcal{M}$ in Fig. 6. In Figs. 6(a) and 6(b), the variations of the rational number $q$ with particle energy are shown. It is found that $q$ increases slowly with energy $E$ initially and then increases rapidly after reaching the maximum energy. In Fig. 6(a), the variation of $q$ with $\epsilon$ is shown for the case where the hair parameter $\ell/\mathcal{M}=1$ and $\epsilon$ takes different values. It is observed that $\epsilon$ has a significant impact on the distribution of qEq-E. In Fig. 6(b), we fix $\epsilon=0.5$ and explore the influence of different $\ell/\mathcal{M}$ on $q$. The results show that the closer the system is to the Schwarzschild case, the larger the maximum energy corresponding to $q$. On the other hand, in Figs. 6(c) and 6(d), the behaviors of $q$ as a function of the $L/\mathcal{M}$ are shown. It is found that $q$ decreases as the orbital angular momentum increases.